OpenStudy (anonymous):

factorise, x^2 -7x +12

OpenStudy (anonymous):

(x-4)(x-3) http://www.wolframalpha.com/input/?i=x%5E2+-7x+%2B12

OpenStudy (apoorvk):

x^2-7x+12 =x^2-3x-4x+12 =x(x-3)-4(x-3) =(x-3)(x-4)

OpenStudy (anonymous):

oh, now I get it. ok, thanks.